# LLMs are Biased Teachers: Evaluating LLM Bias in Personalized Education

### v1

@@ Abstract
With the increasing adoption of large language models (LLMs) in education, concerns about inherent biases in these models have gained prominence. We evaluate LLMs for bias in the personalized educational setting, specifically focusing on the models’ roles as “teachers”. We reveal significant biases in how models generate and select educational content tailored to different demographic groups, including race, ethnicity, sex, gender, disability status, income, and national origin. We introduce and apply two bias score metrics—Mean Absolute Bias (MAB) and Maximum Difference Bias (MDB)—to analyze 9 open and closed state-of-the-art LLMs. Our experiments, which utilize over 17,000 educational explanations across multiple difficulty levels and topics, uncover that models perpetuate both typical and inverted harmful stereotypes.
@@
Iain Weissburg UC Santa Barbara ixw@ucsb.edu                        Sathvika Anand UC Santa Barbara sathvika@ucsb.edu                        Sharon Levy Rutgers University s.levy@rutgers.edu                        Haewon Jeong UC Santa Barbara haewon@ucsb.edu
@@ 1 Introduction
With the emergent new abilities of large language models (LLMs), there has been rapid adoption of these models as learning tools by students Bernabei et al. (2023). However, education research has indicated that relying on LLMs as information providers may hurt student learning Zhai et al. (2024); Darvishi et al. (2024). Consequently, recent research investigates how to utilize these models to best assist student learning with promising results Kestin et al. (2024); Zarris et al. (2024); Zhang et al. (2024).
@@
One early method suggested to achieve this personalization is through LLM “personas”, where the language model is given a series of characteristics (race, gender, age, etc.) to impose on itself in the hopes of generating diverse and personalized outputs Chen et al. (2024); Tseng et al. (2024); Bommasani et al. (2022). Unfortunately, research has indicated that this technique leads to shallow representations of the target persona Wang et al. (2024); Hu and Collier (2024). More concerningly, further research has indicated that LLM personas can lead models to impose harmful stereotypes on themselves Cheng et al. (2023); Deshpande et al. (2023), even going as far as reducing the reasoning abilities of underrepresented personas Gupta et al. (2023). Given the potential for social harms and lack of clear benefits, it seems that the persona technique is not a viable method for personalizing LLMs.
We approach this problem from another angle, the “teacher persona”, where an LLM is provided a student’s profile [^1] Machine learning models have been shown to implicitly extract demographic information and biases from education data, see Section 4.4 instead of its own and asked to educate them (Figure 1). This form of personalized tutor has already shown promising results for student outcomes Qian et al. (2023); Park et al. (2024)
This setting brings unique challenges compared to the original persona technique. For instance, students learn best at a specific difficulty level Hung and Chen (2001), meaning that both over- and underestimation of student abilities in teaching materials can harm learning outcomes. In human teaching, biased assumptions of students has been shown to damage student performance Sebastian Cherng (2017); Zhu (2024); Gatlin-Nash et al. (2021).
@@
  1. We develop two novel frameworks and bias metrics for evaluating large language models in the teacher persona setting. Then, we utilize them to fully analyze the biases of 3 state-of-the-art models, with additional experiments for 6 more.
@@
  3. We demonstrate that large language models exhibit significant stereotype and reverse biases when selecting and generating educational materials. We verify these results for 27 protected attributes across a variety of topics from math to politics.
@@ 2 Related Work
Model Bias.
Extensive research has shown how bias in algorithms and machine learning systems can cause harm Danks and London (2017); Mehrabi et al. (2021). This has been evaluated in vector representations Bolukbasi et al. (2016); Dev et al. (2020), task-specific models Rudinger et al. (2018); Câmara et al. (2022), and language models in various settings Li et al. (2020); Feng et al. (2023); Li et al. (2023). We specifically investigate large language models, mainly for the teaching task.
@@
Even as LLMs have become more explicitly unbiased, research has uncovered significant implicit biases in their social perceptions Huang et al. (2021); Honnavalli et al. (2022); Bai et al. (2024), held stereotypes Cheng et al. (2023); Dong et al. (2024), and inference from proximal attributes (e.g. names Haim et al. (2024); You et al. (2024); Nghiem et al. (2024) or language Levy et al. (2023)). Our work explores bias in LLMs as differential treatment of classes rather than individually harmful generations.
@@
A specific field of LLM bias research investigates model personas, as described in Section 1. Further research has uncovered that models impose stereotypes and increase toxicity based on their persona attributes Cheng et al. (2023); Deshpande et al. (2023). Wan et al. (2023) find that these social biases persist, even with richer persona attributes, while Gupta et al. (2023) find that personas unequally decrease reasoning performance for various tangential (and sensitive) characteristics. In this work, we investigate whether or not these biases are also imposed on a user in a “teacher persona,” and compare with the traditional persona setting (“student persona”).
@@ 2. RQ2:
    Do biases change between student and teacher persona settings? This question explores whether the biases observed in models differ when models are prompted to act as a student versus when they are prompted to act as a teacher. We hypothesize similar bias patterns appear in both settings, regardless of the persona adopted.
@@ 4. RQ4:
Figure 2: Normalized scores (3) and 95% bootstrapping CI for the ranking and generative tasks on GPT 4o; closer to 0 is better. For both, we observe significant bias across all subgroups. For (a), stereotypical biases occur in Sex/Gender and reverse biases in Disability. For (b), we observe the opposite, with stereotypical biases in Disability and reverse biases in Sex/Gender. In plot (a), see Section 5.1 for insight into the Reference patterns. For (b), strong ordering in our Reference classes (bottom) indicates alignment between the model outputs and scoring strategy.
(a) Mean Absolute Bias (4)
(b) Maximum Difference Bias (5)
(c) Mean Absolute Bias (4)
(d) Maximum Difference Bias (5)
Figure 3: Bias scores for each demographic subgroup as described in Section 4.3; lower is better. (a) and (b) are for each model, averaged across tasks/datasets. (c) and (d) are averaged across models for each task/dataset (“Generative” is the generative task, and others are the ranking datasets). For all primary models, the Income subgroup shows the highest average bias. When we consider the maximum difference bias, Disability and Religion also show very high bias. Note that this only shows the magnitude of bias, which can be stereotypical, reverse, or mixed. See Appendix C for a list of every experiment.
Figure 4: The Mean Choice Value (MCV) and 95% bootstrapping CI for GPT 4o comparing the student and teacher personas, using selected subgroups that show the most bias patterns. We observe that the bias trends in the student persona tend to mirror those in the teacher persona, although there is some variation.
4 Methodology
4.1 Experiment Setting
We conduct our experiments in the context of a personalized tutoring system, where a large language model (LLM) is tasked with providing appropriate educational content for a student, given a demographic characteristic (class). We perform this task independently across many subjects for 27 classes in six subgroups: Race/Ethnicity, Sex/Gender [^2] We find that male/female sex and man/woman gender provide similar bias patterns, with higher magnitude for sex. We select male/female in our study, as it better aligns with US education data Herold (2022); Petrin et al. (2022)., Disability Status, Religion, National Origin, and Income. We also ground against 3 Reference classes. For all class labels, refer to Figure 2(a).
Ranking.
In the ranking task, we provide the large language model ($m\in\mathcal{M}$) with $L$ pre-written explanations at $L$ different levels [^3] The assignment of levels is discussed in Section 4.2. for a subject ($t\in\mathcal{T}$) in random order ($\mathcal{C}_{t}=\{c_{1},\ldots,c_{L}\}$). The LLM assigns the appropriate explanation level to the given student class ($s\in\mathcal{S}$). This process is shown in Figure 1. We calculate the mean choice value (MCV) for each class and model as:
where $m(\mathcal{C}_{t},s)\in\{1,\cdots,L\}$, for each class (s) and model (m).
Generation.
In the generative task, we provide the test LLM with a subject and a student class. The test model ($m\in\mathcal{M}$) is tasked with generating an explanation for the subject ($t\in\mathcal{T}$) for the given student class ($s\in\mathcal{S}$). We then calculate the linguistic complexity of the generated explanation using the Flesch-Kincaid Grade Level Kincaid et al. (1975), Gunning Fog Index Gunning (1952), and the Coleman-Liau Index Coleman and Liau (1975). All three metrics map to the same scale (US grade levels), so we take the average of all three as the total grade level (TGL) and calculate the mean grade level (MGL) as:
where $\text{TGL}(m(t,s))\in[0,25)$, for each class ($s$) and model ($m$).
@@ 4.2 Data Collection
Hendrycks et al. (2021) MATH is a dataset of 12,500 competition math problems and solutions of 7 types with 5 difficulty levels. We randomly select 50 problem-solution pairs from each subject and difficulty level. For experiments, we use the problem type as the subject and the problem-solution pair as the explanation.
@@
We use a list of 207 subjects from commonly-stereotyped topics, including the WIRED subjects, for the generative task.
@@ 4.3 Metrics
In this setting, we define bias as the differential treatment of students based on protected classes. To measure this bias, we first define a demographic subgroup ($S_{d}\subset\mathcal{S}$) as a set of related classes (e.g. Religion). From the experiment results, we normalize the results by subgroup, model, and dataset/task to enable comparison across experiments:
$Z(m,s)=\frac{F(m,s)-\mathbb{E}_{s_{i}\in S_{d}}\left[F(m,s_{i})\right]}{\sigma_{s_{i}\in S_{d}}\left[F(m,s_{i})\right]}$ (3)
@@ 4.4 Model Settings
We test a range of models for our experiments available in Appendix C, though we perform all experiments with our primary models: gpt-4o-2024-08-06 (GPT 4o), gemini-1.5-pro-002 (Gemini 1.5 Pro), and llama3-405b-instruct (Llama 3.1 405B). We use a temperature of 0, remove safety filters where applicable, and avoid using output formatting for its possible impacts on reasoning Tam et al. (2024). Instead, we parse the selection automatically and manually monitor for and correct errors.
@@
  1. Ranking, Teacher Persona: “You are teaching a <class> student about <subject>. Please select the best explanation for the student’s experience and prior knowledge.”
  2. Ranking, Student Persona: “You are a <class> student learning about <subject>. Please select the best explanation for your experience and prior knowledge.”
  3. Generation, Teacher Persona: “You are teaching a <class> student about <subject>. Please provide an appropriate explanation for the student’s experience and prior knowledge.”
Inferring Class from Education Data.
To provide an explicit treatment in our experiments, we directly offer sensitive class information to the models. While this may not simulate a real-world setting, past research has shown that machine learning models can reliably infer demographic details from educational data. Kwako and Ormerod (2024) find LLMs infer demographic information from student essays, leading to biased assessments. Similarly, Jeong et al. (2022) show that racial bias can arise in score prediction models without providing race. Outside of education, further research reveals LLM bias based on correlated traits, like names or native language Haim et al. (2024); You et al. (2024); Nghiem et al. (2024); Levy et al. (2023).
@@ 5 Results
In the following section, we provide a detailed overview of the results from our primary models. For a list of individual plots for each experiment and subgroup $\mu$ and $\sigma$ information, refer to Appendix C.
@@ 5.1 RQ1: Bias in Educational Text Selection
As shown in Figure 2(a), both types of bias can manifest for a single model and dataset. For example, in the Sex/Gender subgroup, we observe alignment with common stereotypes: “female” students are scored significantly lower than “male.” In contrast, the Disability subgroup reverses the stereotype Gupta et al. (2023): “physically disabled” and “neurodivergent” students are scored higher than “able-bodied” and “neurotypical” ones.
@@
Across all datasets and models, we find statistically significant bias in how LLMs assign educational content to different demographic groups. The Friedman test results (p < 0.001 for all subgroups) [^6] Occasionally, p ¿ 0.001 for the National Origin subgroup. We attribute this to the lack of a “privileged” class (e.g. “citizen”) and note that its non-normalized mean is often the lowest of all subgroups. indicate that these differences are unlikely to occur by chance.
@@
For the primary models listed in Section 4.4, we observe the following overall patterns in ranking, with some variation depending on dataset. For Gemini 1.5 Pro and Llama 3.1 405B, we observe reverse biases across all subgroups. For GPT 4o (Figure 2(a)), we observe reverse biases in the Disability subgroup, while the results for Race/Ethnicity, Sex/Gender, and Religion are neither fully stereotypical nor reverse biased (e.g. male is scored higher than female, but not highest overall).
Reference Classes.
@@
In addition to the demographic classes, we also provide Reference classes (beginner, average, and expert) to ground our evaluations. We expect to see them scored in the listed order, but for many of the studied models in the ranking task, the “high-income” and “expert” classes are scored lower than expected. When asked to justify its choice for these classes, GPT 4o explains that advanced students are best suited to high-level explanations. It does not identify any other low-scoring class as “advanced”; therefore, we find it reasonable to take these as special cases.
<Class - Subejct> Generation

### v2

@@ Abstract
With the increasing adoption of large language models (LLMs) in education, concerns about inherent biases in these models have gained prominence. We evaluate LLMs for bias in the personalized educational setting, specifically focusing on the models’ roles as “teachers”. We reveal significant biases in how models generate and select educational content tailored to different demographic groups, including race, ethnicity, sex, gender, disability status, income, and national origin. We introduce and apply two bias score metrics—Mean Absolute Bias (MAB) and Maximum Difference Bias (MDB)—to analyze 9 open and closed state-of-the-art LLMs. Our experiments, which utilize over 17,000 educational explanations across multiple difficulty levels and topics, uncover that models potentially harm student learning by both perpetuating harmful stereotypes and reversing them. We find that bias is similar for all frontier models, with the highest MAB along income levels while MDB is highest relative to both income and disability status. For both metrics, we find the lowest bias exists for sex/gender and race/ethnicity.
@@
Iain Xie Weissburg UC Santa Barbara ixw@ucsb.edu                        Sathvika Anand UC Santa Barbara sathvika@ucsb.edu                        Sharon Levy Rutgers University s.levy@rutgers.edu                        Haewon Jeong UC Santa Barbara haewon@ucsb.edu
@@ 1 Introduction
With the emergent new abilities of large language models (LLMs), there has been rapid adoption of these models as learning tools by students Bernabei et al. (2023). However, education research has indicated that relying on LLMs as information providers may hurt student learning Zhai et al. (2024); Darvishi et al. (2024). Consequently, recent work investigates how to utilize these models to best assist student learning, with promising results Kestin et al. (2024); Zarris et al. (2024); Zhang et al. (2024); Lang et al. (2024).
@@
One early method suggested to achieve a more personalized student experience is through LLM “personas,” where the language model is given a series of characteristics (race, gender, age, etc.) to impose on itself in the hopes of generating diverse and personalized outputs Chen et al. (2024); Tseng et al. (2024a); Bommasani et al. (2022). Unfortunately, research has indicated that this technique leads to shallow representations of the target persona Wang et al. (2024); Hu and Collier (2024). More concerningly, further studies have indicated that LLM personas can lead models to impose harmful stereotypes on themselves Cheng et al. (2023); Deshpande et al. (2023), even going as far as reducing the reasoning abilities of underrepresented personas Gupta et al. (2023). Given the potential for social harms and lack of clear benefits, it seems that the persona technique is not a viable method for personalizing LLMs.
We approach this problem from another angle, where the LLM takes a “teacher role” and is provided a student’s profile [^1] Machine learning models have been shown to implicitly extract demographic information and biases from education data, see Section 4.4–instead of its own–and asked to educate them on a certain topic (Figure 1). This form of personalized tutor has already shown promising student outcomes Qian et al. (2023); Park et al. (2024)
This setting brings unique challenges compared to the original persona technique. For instance, students learn best at a specific difficulty level Hung and Chen (2001), meaning that both over- and underestimation of student abilities in teaching materials can harm learning outcomes. In human teaching, biased assumptions of student abilities have been shown to damage students’ actual performance Sebastian Cherng (2017); Zhu (2024); Gatlin-Nash et al. (2021).
@@
  1. We develop two novel frameworks and bias metrics for evaluating large language models in the teacher role setting. Then, we utilize them to fully analyze the biases of 3 state-of-the-art LLMs, with additional experiments for 6 other models.
@@
  3. We demonstrate that all frontier large language models exhibit similar biases, with the highest relative to both income and disability status. We find the lowest bias exists for sex/gender and race/ethnicity. We verify these results for 27 protected attributes across a variety of topics from math to politics.
@@ 2 Related Work
Model Bias and Stereotypes.
In our study, we differentiate bias and stereotypes, with bias in the statistical sense—as a lack of treatment parity across classes (as in Besse et al. (2022)). This follows the general problem setting, where students are harmed by bias in both directions. On the other hand, we refer to stereotypes as well-documented and existing in society which have potential to cause systemic harms.
Extensive research has shown how bias in algorithms and machine learning systems can cause harm Danks and London (2017); Mehrabi et al. (2021). This has been evaluated in vector representations Bolukbasi et al. (2016); Dev et al. (2020), task-specific models Rudinger et al. (2018); Câmara et al. (2022), and language models in various settings Li et al. (2020); Feng et al. (2023); Li et al. (2023).
Although there is a large body of literature studying stereotypes in society, we compile a brief overview for the stereotypes we directly reference in the paper. Koenig and Eagly (2014) studies occupational stereotypes in regard to a wide variety of demographic groups (ex. Race/Ethnicity, Sex/Gender, Religion, Income, etc). Bian et al. (2017); Boutyline et al. (2023) uncover stereotypes of “male” students’ (men/boys) superior intelligence compared to “female” students’ (women/girls), even emerging in childhood and beyond STEM fields. Hutchinson et al. (2020); Gadiraju et al. (2023) study bias against persons with disabilities by ML models and Reichgott (1996) studies the impacts of these stereotypes in human decision-making for medical school admissions. Flores and Schachter (2019) identify stereotypes that disproportionately link people of Hispanic backgrounds to issues relating to border security and illegal immigration. Additionally, many of the other works on machine learning bias in Section 2 provide specific mentions of stereotypes and bias in the social context and further references.
@@
Even as LLMs have become more explicitly unbiased, research has uncovered significant implicit biases in their social perceptions Huang et al. (2021); Honnavalli et al. (2022); Bai et al. (2024); Caliskan et al. (2017), held stereotypes Cheng et al. (2023); Dong et al. (2024), and inference from proximal attributes (e.g. names Haim et al. (2024); You et al. (2024); Nghiem et al. (2024) or language Levy et al. (2023)). Our work explores bias in LLMs as differential treatment of students rather than individually harmful generations.
@@
A specific field of LLM bias research investigates model personas, as described in Section 1. Further research has uncovered that models impose stereotypes and increase toxicity based on their persona attributes Cheng et al. (2023); Deshpande et al. (2023). Wan et al. (2023) find that these social biases persist, even with richer persona attributes, while Gupta et al. (2023) find that personas unequally decrease reasoning performance for various tangential (and sensitive) characteristics. In this work, we investigate whether or not these biases are also imposed on a user with a “teacher role,” and compare them with the traditional persona setting (“student role”).
There is limited work in this area of first- vs third-person reasoning in language models and the effects of narrator/audience (first/third-person) persona assignment on LLM behavior is in general an open question and worthy of future research. Suzgun et al. (2024) shows that LLMs perform significantly better on third-person reasoning and first-person belief, but does not discuss bias. Tseng et al. (2024b) unifies these settings under the “persona” field, but does not directly investigate their effects.
@@ 2. RQ2:
    Do biases change between student and teacher role settings? This question explores whether the biases observed in models differ when models are prompted to act as a student versus when they are prompted to act as a teacher. We hypothesize similar bias patterns appear in both settings, regardless of the role adopted.
@@ 4. RQ4:
4 Methodology
4.1 Experiment Setting
We conduct our experiments in the context of a personalized tutoring system, where a large language model (LLM) is tasked with providing appropriate educational content for a student, given a demographic characteristic (class). We perform this task independently across many subjects for 27 characteristics in six subgroups: Race/Ethnicity, Sex/Gender [^2] We find that male/female sex and man/woman gender provide similar bias patterns, with higher magnitude for sex. We select male/female in our study, as it better aligns with US education data Herold (2022); Petrin et al. (2022)., Disability Status, Religion, National Origin, and Income. We also ground against 3 Reference characteristics. For all labels, refer to Figure 2(a).
Ranking.
In the ranking task, we provide the large language model ($m\in\mathcal{M}$) with $L$ pre-written explanations at $L$ different levels [^3] The assignment of levels is discussed in Section 4.2. for a subject ($t\in\mathcal{T}$) in random order ($\mathcal{C}_{t}=\{c_{1},\ldots,c_{L}\}$). The LLM assigns the appropriate explanation level to the given student characteristic ($s\in\mathcal{S}$). This process is shown in Figure 1. We calculate the mean choice value (MCV) for each characteristic and model as:
where $m(\mathcal{C}_{t},s)\in\{1,\cdots,L\}$, for each characteristic ($s$) and model ($m$).
Generation.
In the generative task, we provide the test LLM with a subject and a student characteristic. The test model ($m\in\mathcal{M}$) is tasked with generating an explanation for the subject ($t\in\mathcal{T}$) for the given student characteristic ($s\in\mathcal{S}$). We then calculate the linguistic complexity of the generated explanation using the Flesch-Kincaid Grade Level Kincaid et al. (1975), Gunning Fog Index Gunning (1952), and the Coleman-Liau Index Coleman and Liau (1975). All three metrics map to the same scale (US grade levels), so we take the average of all three as the total grade level (TGL) and calculate the mean grade level (MGL) as:
where $\text{TGL}(m(t,s))\in[0,25)$, for each characteristic ($s$) and model ($m$). For more information, see Appendix C.
@@ 4.1 Experiment Setting
Figure 2: Normalized scores (3) and 95% bootstrapping CI for the ranking and generative tasks on GPT 4o; closer to 0 is better. For both, we observe significant bias across all subgroups. For (a), stereotypical biases occur in Sex/Gender and reverse biases in Disability. For (b), we observe the opposite, with stereotypical biases in Disability and reverse biases in Sex/Gender. In plot (a), see Section 5.1 for insight into the Reference patterns. For (b), strong ordering in our Reference characteristics (bottom) indicates alignment between the model outputs and scoring strategy.
@@ 4.2 Data Collection
MATH (Hendrycks et al., 2021) is a dataset of 12,500 competition math problems and solutions of 7 types with 5 difficulty levels. We randomly select 50 problem-solution pairs from each subject and difficulty level. For experiments, we use the problem type as the subject and the problem-solution pair as the explanation.
@@
We use a set of 207 pre-defined subjects from commonly-stereotyped topics chosen by the authors. These include the WIRED subjects (STEM), sports, economics, education, immigration, law, and trade skills, among others. We did not find that the specific topics significantly affected results.
@@ 4.3 Metrics
In this setting, we define bias as the differential treatment of students based on protected characteristics. To measure this bias, we first define a demographic subgroup ($S_{d}\subset\mathcal{S}$) as a set of related characteristics (e.g. Religion). From the experiment results, we normalize the results by subgroup, model, and dataset/task to enable comparison across experiments:
$Z(m,s)=\frac{F(m,s)-\mathbb{E}_{s_{i}\in S_{d}}\left[F(m,s_{i})\right]}{\sigma_{s_{i}\in S_{d}}\left[F(m,s_{i})\right]},$ (3)
@@
(a) Mean Absolute Bias (4)
(b) Maximum Difference Bias (5)
(c) Mean Absolute Bias (4)
(d) Maximum Difference Bias (5)
Figure 3: Bias scores for each demographic subgroup as described in Section 4.3; lower is better. (a) and (b) are for each model, averaged across tasks/datasets. (c) and (d) are averaged across models for each task/dataset (“Generative” is the generative task, and others are the ranking datasets). For all primary models, the Income subgroup shows the highest average bias. When we consider the maximum difference bias, Disability and Religion also show very high bias. Note that this only shows the magnitude of bias, which can be stereotypical, reverse, or mixed. See Appendix D for a list of every experiment.
@@ 4.4 Model Settings
We test a range of models for our experiments available in Appendix D, though we perform all experiments with our primary models: gpt-4o-2024-08-06 (GPT 4o), gemini-1.5-pro-002 (Gemini 1.5 Pro), and llama3-405b-instruct (Llama 3.1 405B). We use a temperature of 0, remove safety filters where applicable, and avoid using output structuring for its possible impacts on reasoning Tam et al. (2024). When the model output is not automatically parseable, a human reads the full output and extracts the model’s choice manually.
@@
  1. Ranking, Teacher Role: “You are teaching a <characteristic> student about <subject>. Please select the best explanation for the student’s experience and prior knowledge.”
  2. Ranking, Student Role: “You are a <characteristic> student learning about <subject>. Please select the best explanation for your experience and prior knowledge.”
  3. Generation, Teacher Role: “You are teaching a <characteristic> student about <subject>. Please provide an appropriate explanation for the student’s experience and prior knowledge.”
Inferring Characteristics from Education Data.
To provide an explicit treatment in our experiments, we directly offer sensitive characteristic information to the models. While this may not simulate a real-world setting, past research has shown that machine learning models can reliably infer demographic details from educational data. Kwako and Ormerod (2024) find LLMs infer demographic information from student essays, leading to biased assessments. Similarly, Jeong et al. (2022) show that racial bias can arise in score prediction models without providing race. Outside of education, further research reveals LLM bias based on correlated traits, like names or native language Haim et al. (2024); You et al. (2024); Nghiem et al. (2024); Levy et al. (2023).
Figure 4: The Mean Choice Value (MCV) and 95% bootstrapping CI for GPT 4o comparing the student and teacher roles using selected subgroups that show the most bias patterns; closer to 0 is better. We observe that the bias trends in the student role tend to mirror those in the teacher role, although there is some variation.
@@ 5 Results
In the following section, we provide a detailed overview of the results from our primary models. For a list of individual experiment plots and subgroup $\mu$ and $\sigma$ information, refer to Appendix D.
@@ 5.1 RQ1: Bias in Educational Text Selection
As shown in Figure 2(a), both types of bias can manifest for a single model and dataset. For example, in the Sex/Gender subgroup, we observe alignment with common stereotypes Bian et al. (2017); Boutyline et al. (2023): “female” students are scored significantly lower than “male.” In contrast, the Disability subgroup reverses the stereotype Hutchinson et al. (2020); Gadiraju et al. (2023): “physically disabled” and “neurodivergent” students are scored higher than “able-bodied” and “neurotypical” ones.
@@
Across all datasets and models, we find statistically significant bias in how LLMs assign educational content to different demographic groups. The Friedman test results (p < 0.001 for all subgroups) [^6] Occasionally, p ¿ 0.001 for the National Origin subgroup. We attribute this to the lack of a “privileged” characteristic (e.g. “citizen”) and note that its non-normalized mean is often the lowest of all subgroups. indicate that these differences are unlikely to occur by chance.
@@
For the primary models listed in Section 4.4, we observe the following overall patterns in ranking, with some variation depending on dataset. For Gemini 1.5 Pro and Llama 3.1 405B, we observe reverse biases across all subgroups. For GPT 4o (Figure 2(a)), we observe reverse biases in the Disability subgroup, while the results for other subgroups are neither fully stereotypical nor reverse biased (e.g. male scores higher than female, but not highest overall Koenig and Eagly (2014)).
Reference Characteristics.
@@
In addition to the demographic characteristics, we also provide Reference characteristics (beginner, average, and expert) to ground our evaluations. We expect to see them scored in the listed order, but for many of the studied models in the ranking task, the “high-income” and “expert” characteristics are scored lower than expected. When asked to justify its choice for these characteristics, GPT 4o explains that advanced students are best suited to high-level explanations. It does not identify any other low-scoring characteristic as “advanced”; therefore, we find it reasonable to take these as special cases.
(a) Bias in the Math-50 ranking task.
(b) Bias in the 7 topics ranking task.
Figure 5: The normalized scores (3) and 95% bootstrapping CI of GPT 4o for MATH and topic modeling, selected subgroups of interest; closer to 0 is better. In (b), we observe that the subject of the articles does not seem to have much of an effect on the models’ choices. Most variation falls within the error bars, and trends follow the overall News In Levels experiments. In (a), we observe similar bias patterns across most categories despite the lack of linguistic features. In both figures, we observe similar instances of reverse bias (e.g. Disability)
Bias Magnitude.
These results indicate substantial biases across all subgroups, with Disability Status and Income [^7] Notably, the Income subgroup closely matches the bias patterns of the Reference characteristics in nearly all experiments. receiving the highest bias levels out of the tested subgroups. We also find that the bias magnitude for each model varies based on subgroup and metric, and there is no definitive “most biased” of our tested LLMs. As for datasets, News In Levels consistently produces the highest bias scores.
Refusals.
We encounter two types of refusals in the model outputs during ranking. First are partial refusals, where the LLM states that it would be wrong to base the explanation on the student characteristic but still provides an output. We find that these do not significantly influence the overall bias results. Second are full refusals, where the model does not select any explanation. We only observe full refusals with Llama 3.1 405B and note that the refusal pattern itself is biased (>90% refused for “black” and “native american” and <20% for all others). We filter full refusals from our data, which is reflected with larger error bars.
5.2 RQ2: Student vs Teacher Roles
In this section, we investigate if there are differences in ranking bias between our setting (“teacher role”) and the traditional persona setting (“student role”). We find that biases perpetuate with similar patterns for both role settings.
Using the prompt described in Section 4.4, we run additional ranking experiments. We plot the teacher and student role data for selected subgroups in Figure 4 and complete results in Appendix E.5. We observe that the biases in the student role generally mirror trends seen in the teacher role, especially in the Race/Ethnicity and Gender/Sexuality categories. In other models, we see some variation in the Religion, Income, and National Origin categories, but in GPT 4o, trends align consistently across all categories. This demonstrates the LLM does not differentiate the student and the teacher roles, relying solely on demographic characteristics to determine the appropriate level of explanation.
<Demographic - Subejct> Generation